# Aubry-André Transition for an Active Undulator

### v1

@@ Acknowledgments
Furthermore, because the body does not deform along $x$, the laboratory and body reference frames may be related by $X=X_{0}+x$, where $X_{0}$ is the position of the anterior-most point (the tail) of the undulator along the $X$ axis. Therefore, substituting (S7) into (S6) and noting that $dX_{0}/dt=dX/dt$ yields the following differential equation
@@
$\xi(X)=\Delta\sin(kX)+\Gamma\sin(\beta kX)$ (S9)
the integral on the rhs of S8 may be integrated analytically, producing a non-linear first order ODE that can be solved numerically. For $k=2$, it is written as
$\frac{1}{\pi^{2}y_{0}^{2}}\frac{dX_{0}}{dt}~{}=$ $~{}\Delta\cos{4\pi(X_{0}-t)}~{}+$ (S10)
$~{}\frac{\Gamma}{\beta(\beta^{2}-4)}\times$
$~{}~{}~{}~{}\bigg{(}\beta(\beta+2)~{}\sin{\beta\pi}~{}\cos{\pi(2\beta X_{0}-4t+\beta)}$
$~{}~{}~{}~{}+\beta(\beta-2)~{}\sin{\beta\pi}~{}\cos{\pi(2\beta X_{0}+4t+\beta)}$
$~{}~{}~{}~{}-(\beta+2)(\beta-2)(\sin{2\beta X_{0}}-\sin{2\beta(X_{0}+1)}\bigg{)}.$
To calculate $X_{0}(t)$ we numerically solved (S10) using Mathematica’s numerical ODE solver. We note that for the case where $\Gamma=0$ (a single periodic drag function), (S10) is similar to the non-uniform oscillator equation, with stable fixed points at $X_{0}-t=0$, which correspond with uniform motion.

### v2

@@ 1 Results and Discussion
Furthermore, because the body does not deform along $x$, the laboratory and body reference frames may be related by $X=X_{0}+x$, where $X_{0}$ is the position of the anterior-most point (the tail) of the undulator along the $X$ axis. Therefore, substituting (11) into (10) and noting that $dX_{0}/dt=dX/dt$ yields the following differential equation
@@
$\xi(X)=\Delta\sin(kX)+\Gamma\sin(\beta kX)$ (13)
the integral on the rhs of 12 may be integrated analytically, producing a non-linear first-order ODE that can be solved numerically. For $k=2$, it is written as
$\frac{1}{\pi^{2}y_{0}^{2}}\frac{dX_{0}}{dt}\leavevmode\nobreak\ =$ $\leavevmode\nobreak\ \Delta\cos{4\pi(X_{0}-t)}\leavevmode\nobreak\ +$ (14)
$\leavevmode\nobreak\ \frac{\Gamma}{\beta(\beta^{2}-4)}\times$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bigg{(}\beta(\beta+2)\leavevmode\nobreak\ \sin{\beta\pi}\leavevmode\nobreak\ \cos{\pi(2\beta X_{0}-4t+\beta)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\beta(\beta-2)\leavevmode\nobreak\ \sin{\beta\pi}\leavevmode\nobreak\ \cos{\pi(2\beta X_{0}+4t+\beta)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -(\beta+2)(\beta-2)(\sin{2\beta X_{0}}-\sin{2\beta(X_{0}+1)}\bigg{)}.$
To calculate $X_{0}(t)$ we numerically solved (14) using Mathematica’s numerical ODE solver. We note that for the case where $\Gamma=0$ (a single periodic drag function), (14) is similar to the non-uniform oscillator equation, with stable fixed points at $X_{0}-t=0$, which correspond with uniform motion.
[…]
This work was funded by NSF 2310751 and DOD W911NF2110033.
[…]
[2]